. The prescriber ordered Quinidine gluconate 600 mg TID IM. The drug is available as 200 mg/mL. How many mL (whole number) will you give for each dose? 
Answer #1

ANSWER.

* Calculation of volume(ml) of drug required.

         Formula

             Volume(ml) = Desired dose(D)/Dose in hand(H) * Quantity(Q).

Here

       Desired dose(D) = 600 mg.

       Dose in hand(H) = 200 mg.

           Quantity(Q) = 1 ml

As per above formula

   Volume(ml) of drug required = D/H * Q

                                                  = 600/200 * 1

                                                  = 3 * 1

                                                  = 3 ml per dose.